# Leave No Document Behind: Benchmarking Long-Context LLMs with Extended Multi-Doc QA

### v1

@@ 4.2 Main Results
We assess six advanced LLMs on the Loong benchmark. The main results are shown in Table 3 and Table 4. We can see that Gemini-Pro-1.5 shows the best overall performance, especially excelling in the processing of ultra-long context within $\mathtt{Set3}$ and $\mathtt{Set4}$. Its comprehensive score reached 55.37 with the perfect rate of 27%, followed by GPT-4o. Besides, the long-context modeling capacity of open-source models still falls short when compared to that of the most powerful closed-source models in the Loong. Additionally, larger-parameter models outperform their smaller counterparts within the same window size, indicating the advantages of scaling up model sizes for improved long-context modeling. The overall assessment results highlight that even the most advanced long-context LLMs currently fail to achieve passing marks, particularly in terms of the perfect rate. This suggests that there exists significant room for improvement in the long-context modeling capabilities of LLMs.
@@ 4.5 RAG or Not
Relying on RAG cannot resolve all the problems associated with long-text modeling. To genuinely improve the long-context modeling capability, stronger training methods and effective training on longer texts are required, rather than merely integrating the RAG module.
@@ B.4 Range Awareness
Question: Which company has the highest ’Total Current Liabilities’?
@@ Appendix E Comparison of Evidence Distribution
In the same multi-document question-answering task, we compared the distribution of evidence related to the answers in the context for Loong (Ours) and Longbench Bai et al. (2023b). As shown in Table 9, we present data examples from Loong and LongBench. It can be observed that although Longbench contains a large number of passages, the evidence is only distributed within Passage 1. In contrast, in our Loong, the evidence is distributed across every document, requiring the model to understand each document in order to provide the correct answer.

### v2

@@ 4.2 Main Results
We assess seven advanced LLMs on the Loong benchmark. The main results are shown in Table 3 and Table 4. We can see that Gemini-1.5-pro shows the best overall performance, especially excelling in the processing of ultra-long context within $\mathtt{Set3}$ and $\mathtt{Set4}$. Its comprehensive score reached 55.37 with the perfect rate of 27%, followed by GPT-4o. Besides, the long-context modeling capacity of open-source models still falls short when compared to that of the most powerful closed-source models in the Loong. Additionally, larger-parameter models outperform their smaller counterparts within the same window size, indicating the advantages of scaling up model sizes for improved long-context modeling. The overall assessment results highlight that even the most advanced long-context LLMs currently fail to achieve passing marks, particularly in terms of the perfect rate. This suggests that there exists significant room for improvement in the long-context modeling capabilities of LLMs. We also compare the current results with other current Long-Context benchmarks, which can be seen in Appendix F.
@@ 4.5 RAG or Not
We also provide analytical experiment to show that RAG does not cover all evidence, which can be seen in Appendix G. Relying on RAG cannot resolve all the problems associated with long-text modeling. To genuinely improve the long-context modeling capability, stronger training methods and effective training on longer texts are required, rather than merely integrating the RAG module.
@@ 5 Conclusion
Acknowledgments
Min Yang was supported by National Key Research and Development Program of China (2022YFF0902100), National Natural Science Foundation of China (Grant No. 62376262), the Natural Science Foundation of Guangdong Province of China (2024A1515030166), Shenzhen Science and Technology Innovation Program (KQTD20190929172835662), Shenzhen Basic Research Foundation (JCYJ20210324115614039).
@@ B.4 Range Awareness
Question: How many companies have ’Total Shares Outstanding’ exceeding 10,000,000 shares?
@@ Appendix E Comparison of Evidence Distribution
In the same multi-document question-answering task, we compared the distribution of evidence related to the answers in the context for Loong (Ours) and Longbench Bai et al. (2024). As shown in Table 11, we present data examples from Loong and LongBench. It can be observed that although Longbench contains a large number of passages, the evidence is only distributed within Passage 1. In contrast, in our Loong, the evidence is distributed across every document, requiring the model to understand each document in order to provide the correct answer.
Appendix F Comparison of Results with Other Benchmarks
Due to the rapid iteration of models and the high cost of closed-source model APIs, previous studies have rarely evaluated models that support context windows of 128k or more. Evaluating the latest models with ultra-long context windows and expensive closed-source models is our advantage compared to previous methods. Users can evaluate the gap between their models and the latest open-source and closed-source models at a low cost. To discover more interesting conclusions, we compared Loong with RULER  Hsieh et al. (2024), and the most recent work, NOCHA  Karpinska et al. (2024). The experimental results are shown in the Table 5.
From the result, we discovered that: (1) Although RULER focuses on synthetic tasks and NOCHA focuses on the domain of novels, the performance of the models on the three benchmarks is essentially consistent. (2) Gemini-1.5-pro achieved the best results in Loong and RULER, but GPT-4o had a significant lead on NOCHA. We believe that since NOCHA’s data sources only include novels, GPT-4o might have stronger capabilities in the domain of novel literature. (3) The large-parameter Qwen2-72B-Instruct underperformed in RULER compared to the smaller-parameter models GLM4-9B-Chat and Llama3.1-8B, which is inconsistent with the results of Loong. We believe that RULER emphasizes synthetic tasks, where the performance of such tasks is as sensitive as Needle In A Haystack. Specifically, simply adding some synthetic task data to the training corpus can yield good results. In real-world tasks like Loong, such anomalous phenomena generally do not occur.
Models Loong RULER NOCHA
GPT-4o 53.47 - 55.80
GPT-4-Turbo - 91.60 40.20
Gemini-1.5-pro 55.37 95.80 48.10
Claude3.5-Sonnet 48.85 - 41.00
Qwen2-72B-Instruct 43.29 85.90 43.35
GLM4-9B-Chat 38.31 89.90 27.07
Llama3.1-8B 36.31 88.30 16.53
Phi-3-mini-3.8B 14.54 68.80 9.30
Table 5: Compare Loong with RULER and NOCHA.
Appendix G Results of Recall Rate by RAG
To show that RAG does not cover all evidence, we sampled all the questions from the financial test case in the Loong and analyzed RAG‘s retrieval results. Since the answers to our questions are distributed across all documents on average, we evaluated whether the retrieved top-k passages cover all documents to reflect the RAG’s recall of evidence. This metric will be higher than the actual evidence recall rate because even if all required documents are retrieved, the specific passages may not necessarily contain the evidence. If the model fails to retrieve all documents, it will certainly be unable to retrieve all evidence. The evaluation metric used is Recall@n, where n represents the number of retrieved passages. If the n passages contain all documents, Recall@n is set to 1; otherwise, it is 0. The results can be seen in Table 6. Evidently, in the Loong dataset, RAG struggles to retrieve all relevant documents. Even when the topK reaches 50, the highest recall is only 0.64. In reality, the recall rate for all evidence will be even lower.
Model Recall@10 Recall@20 Recall@30 Recall@40 Recall@50
Openai Embedding, chunk size=1024 0.20 0.34 0.43 0.51 0.58
Openai Embedding, chunk size=2048 0.20 0.36 0.48 0.57 0.62
BGE Embedding, chunk size=1024 0.19 0.34 0.44 0.53 0.59
BGE Embedding, chunk size=2048 0.21 0.38 0.49 0.58 0.64
Table 6: The Recall Rate of all documents by using RAG.
Appendix H Detailed URL of Document Source
China Judge Online:
U.S. SEC:
cninf:
Arxiv:
Semantic Scholar: